Write a function for the line, in the form f(x)=mx+b, using the information given below.
A linear function containing the points in the following table
x
3
5
and f(x)
-2
-8

Answer by scott8148(6628)

substituting values from the table __ -2=m(3)+b __ -2=3m+b

substituting again __ -8=m(5)+b __ -8=5m+b

subtracting the first equation from the second __ -6=2m __ -3=m

substituting __ -2=-3(3)+b __ -2=-9+b __ 7=b